OpenStudy (khally92):

find the first derivative (1-2x)^2(x-3) by expandind the equation.. i got 12x^2-32x+13. is dat the correct ans please

OpenStudy (anonymous):

2(1-2x)(-2)(x-3)+(1-2x)^2=-4(1-2x)(x-3)+(1-2x)^2=-4(x-3-2x^2+6x)+1-4x+4x^2= -4x +12 +8x^2-24x+1-4x+4x^2=12x^2-32x+13

OpenStudy (khally92):

i need to also fyn the second derivative so i was told thatif i expand the whole bracket it will be much easier. please

OpenStudy (khally92):

woah amazing that means i got it ryt?

OpenStudy (anonymous):

yes, you got it

OpenStudy (khally92):

thanks alot. i really appreciate.

OpenStudy (khally92):

hello please can we continue